# Semiclassical measure of the propagation between two topological insulators

### v2

@@ 1.3 Main result
$\frac{\nabla\mathrm{m}(\gamma(s))}{\left|\nabla\mathrm{m}(\gamma(s))\right|}=\begin{pmatrix}\cos\theta(s)\\ \sin\theta(s)\end{pmatrix}.$
We will denote the dual variable of $(s,y)$ by $(\sigma,\eta)$. Let us consider for all $(s,y,\sigma)\in\mathbb{R}^{3}$ and for all $n\in\mathbb{Z}_{\neq 0}$,
$\lambda_{0}(\sigma)\coloneqq-\sigma\;\;,\;\;\lambda_{n}(s,\sigma)\coloneqq\mathrm{sgn}(n)\sqrt{\sigma^{2}+2|n||\nabla\mathrm{m}(\gamma(s))|}.$ (5)
Although the function $\lambda_{0}$ does not depend on $s$, we also will denote it by $\lambda_{0}(s,\sigma)$.
A consequence of our main result (in Section 3.3) is the following asymptotic description of the position density of the solution of equation (1) for initial data which are wave packets concentrating on the curve $\mathcal{C}$.
Theorem 1.4.

### v3

@@ 2.1 Straightening the edge
A consequence of Assumption 1.3 is that for all $\delta_{0}\in\left(0,1\right]$, the function $\Phi_{\delta_{0}}$ is a diffeomorphism from $\mathbb{R}\times\mathrm{I}_{\delta_{0}}$ to
[…]
$\varphi^{\varepsilon}_{t}\coloneqq\mathrm{U}\chi\psi^{\varepsilon}_{t}.$ (2.1)
Lemma 2.2 (Dirac equation in normal geodesic coordinates).
The family $\left(\varphi^{\varepsilon}_{t}\right)_{\varepsilon>0}$ satisfies, on $\mathbb{R}_{t}\times\mathbb{R}_{s}\times\mathrm{I}_{1/2}$, the following equation
$\varepsilon\mathrm{D}_{t}\varphi^{\varepsilon}_{t}+\mathrm{H}^{\mathrm{E}}_{\varepsilon}\varphi^{\varepsilon}_{t}=0.$ (2.2)
Proof.
We apply $\mathrm{U}\chi$ to equation (1.1), obtaining
A direct computation gives
[…]
$\Big(\mathrm{U}\left[\mathrm{H}_{\varepsilon},\chi\right]\psi_{t}^{\varepsilon}\Big)(s,y)=0,\quad(t,s,y)\in\mathbb{R}_{t}\times\mathbb{R}_{s}\times\mathrm{I}_{1/2}.$
Pull-back of semiclassical measure.
We establish the connection between the semiclassical measure of the family $\left(\psi_{t}^{\varepsilon}\right)_{\varepsilon>0}$ and the two-scale semiclassical measure of the family $\left(\varphi_{t}^{\varepsilon}\right)_{\varepsilon>0}$ defined in (2.1) via the change of variables introduced in Proposition 2.1.
Lemma 2.3 (Semiclassical measure and normal geodesic coordinates).
All semiclassical measures $\mu_{t}$ of the family $(\psi^{\varepsilon}_{t})_{\varepsilon>0}$ associated with the sequence $(\varepsilon_{k})_{k\in\mathbb{N}}$ satisfy
where $\rho_{t}$ is the semiclassical measure of the family $\left(\varphi_{t}^{\varepsilon}\right)_{\varepsilon>0}$ on $\mathbb{R}_{s}\times\mathrm{I}_{1/2}\times\mathbb{R}_{\sigma,\eta}^{2}\times\mathbb{R}_{t}$ associated with the sequence $\left(\varepsilon_{k}\right)_{k\in\mathbb{N}}$.
Proof.
Let $a\in\mathscr{C}_{c}^{\infty}\left(\mathbb{R}^{4},\mathbb{C}^{2,2}\right)$ with $\mathrm{supp}(a)\subset\Omega_{1/2}\times\mathbb{R}^{2}$. Then, for all $N\in\mathbb{N}$, we have
Since $\rm U=U_{\theta(s)}V_{\Phi}$, with $\rm U_{\theta(s)}$ depending only on $s$ and $\rm V_{\Phi}$ a change of variables, standard results in semiclassical pseudodifferential calculus (see (Zworski, 2012, Theorem 9.3)) give
$+\mathcal{O}\left(\varepsilon\right).$
Assuming the sequence $\left(\varepsilon\right)_{\varepsilon>0}$ realizes the semiclassical measure $\mu_{t}$ of $(\psi^{\varepsilon}_{t})_{\varepsilon>0}$, passing to the limit $\varepsilon\to 0$ gives
$=\int_{\mathbb{R}\times\mathrm{I}_{1/2}\times\mathbb{R}^{2}}\mathrm{Tr}\left(a\left(\Phi(s,y),{}^{t}\nabla\left(\Phi^{-1}\right)\Big(\Phi(s,y)\Big)\left[\sigma,\eta\right]\right)\Big(\mathrm{U}_{\theta(s)}^{\dagger}\mathrm{d}\rho_{t}(s,y,\sigma,\eta)\mathrm{U}_{\theta(s)}\Big)\right)$
Finally, restricting the measure $\rho_{t}$ to the curve $\mathcal{C}$ gives the semiclassical measure associated with $(\psi^{\varepsilon}_{t})_{\varepsilon>0}$ near $\mathcal{C}$. ∎
2.2 Two-scale analysis of concentrations
In this section, we will consider two-scale semiclassical observables and quantization defined in (1.11) at the scale $\sqrt{\varepsilon}$. We first remark that for all $a\in\mathcal{A}$,
$\mathbf{op}^{\mathcal{C}}_{\varepsilon}(a)$ $=\mathrm{Op}_{\varepsilon}\left(a\left(s,y,\sigma,\eta,\frac{(y,\sigma,\eta)}{\sqrt{\varepsilon}}\right)\right)$
$=\Lambda_{\sqrt{\varepsilon}}^{\dagger}\mathrm{Op}_{1}\Big(a\left(s,\sqrt{\varepsilon}y,\varepsilon\sigma,\sqrt{\varepsilon}\eta,y,\sqrt{\varepsilon}\sigma,\eta\right)\Big)\Lambda_{\sqrt{\varepsilon}},\ \left(\text{semiclassical in }\sqrt{\varepsilon}\right)$ (2.3)
with